# Learning Reward Machines from Partially Observed Optimal Policies

### v1

@@ II-D Problem Statement
    Does there always exist a depth $l^{*}$ such that, given the labeled MDP model $\mathcal{M}_{L}$, a bound $u_{\mathrm{max}}$ on the number of nodes of the underlying reward machine, and the depth-$l^{*}$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l^{*}}$ of the true prefix tree policy, it is possible to learn a reward machine that is policy-equivalent to the underlying one?
@@
In what follows, we first provide an algorithm that takes the labeled MDP model $\mathcal{M}_{L}$, the depth-$l$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l}$ of the prefix tree policy $\pi_{\mathrm{PTP}}^{\mathrm{true}}$ for some arbitrary $l$, and the bound $u_{\mathrm{max}}$, and computes a reward machine that induces a prefix tree policy $\pi_{\mathrm{PTP}}^{\mathrm{learned}}$ with the same depth-$l$ restriction, i.e., $\pi_{\mathrm{PTP}}^{\mathrm{learned},l}=\pi_{\mathrm{PTP}}^{\mathrm{true},l}$. Then, we show the existence of a sufficient depth $l^{*}$ in (P1), for which this algorithm solves problem (P2).
III Methodology
III-A SAT Encoding
@@ III-A1 Determinism
III-A2 Full Specification
This constraint, also known as being input-enabled [28], ensures that all labels generate valid transitions at all states. The corresponding Boolean constraints are:
$\forall i,\forall k,\exists j\text{ such that }b_{ijk}=1.$ (7)
We can combine the conditions of Sections III-A1 and III-A2 into one condition on each $B_{k}$ enforcing that each row has exactly one entry with value $1$.
III-A3 Negative Examples
@@
Proof:
It follows from Eq. (4). ∎
@@
$\mathcal{E}^{-}_{l}=\{(\sigma,\sigma^{\prime})\mid\pi_{\mathrm{PTP}}^{l}(a|s,\sigma)\neq\pi_{\mathrm{PTP}}^{l}(a|s,\sigma^{\prime})\text{ for some }s,a\}$
Let $\sigma=k_{1}k_{2}\cdots k_{l}$ and $\sigma^{\prime}=k_{1}^{\prime}k_{2}^{\prime}\cdots k_{m}^{\prime}$ be two propositional prefixes that lead to different policies in the same state, therefore $(\sigma,\sigma^{\prime})\in\mathcal{E}^{-}_{l}$. We encode the condition given by Lemma 1 into Boolean constraints as:
$(B_{k_{l}}^{\intercal}B_{k_{l-1}}^{\intercal}\cdots B_{k_{1}}^{\intercal}e_{1})\bigwedge$ $(B_{k_{m}^{\prime}}^{\intercal}B_{k_{m-1}^{\prime}}^{\intercal}\cdots B_{k_{1}^{\prime}}^{\intercal}e_{1})$
$=$ $\begin{bmatrix}0&0&\cdots&0\end{bmatrix}^{\intercal},$ (8)
where $\bigwedge$ is the element-wise AND of the matrices and $e_{1}\triangleq[1,0,\cdots,0]^{\intercal}$ indicates that the paths start from the initial node. Our algorithm adds the Boolean constraint given in Eq. (III-A3) for each element of $\mathcal{E}^{-}$. The significance of encoding negative examples is that it eliminates the learning of trivial reward machines. In particular, our method never learns a trivial one-state reward machine with all self-transitions [9] as long as there is at least one negative example in our prefix tree policy.
III-A4 Non-Stuttering
A reward machine is said to be non-stuttering if when a proposition transitions into a node, that same proposition can not transition out of the node, i.e., for all $(u,a,u^{\prime})\in\mathcal{U}\times\mathrm{AP}\times\mathcal{U}$: $\delta_{\mathbf{u}}(u,a)=u^{\prime}\implies\delta_{\mathbf{u}}(u^{\prime},a)=u^{\prime}$. This is related to multi-stage tasks where the particular duration spent on a subtask (i.e., satisfying a given atomic proposition) is not important [29]. When the reward machine is a priori known to be non-stuttering, this extra condition can be included in the SAT problem. The main significance of this condition is trace-compression [9], by which we can reduce the number of negative examples in $\mathcal{E}^{-}$ by only keeping the shortest negative examples among the equivalent ones. In this case, two negative examples are equivalent if between two pairs, the corresponding label sequences differ only by the same proposition repeated consecutively more than one time. We encode non-stuttering into boolean constraints as follows:
@@ III-A4 Non-Stuttering
III-B Algorithm
@@
The last step is finding the rewards that render the product policy $\pi_{\mathrm{Prod}}^{\mathrm{learned}}$ optimal for the product MDP model $(\mathcal{M}/r)^{\mathsf{learned}}$. This is a standard IRL problem without reward machines where the special structure of the rewards on the product can be represented as features. We solve this step using the method developed in [26]. Featurization (see, [26, Section 4]) is used to enforce that the reward function of the product can be written as $r((s,u),a,(s^{\prime},u^{\prime}))=\delta_{\mathbf{r}}(u,L(s^{\prime}))$, with $u^{\prime}=\delta_{\mathbf{u}}(u,L(s^{\prime}))$. This IRL method gives us the corresponding reward function $\delta_{\mathbf{r}}$.
@@
$n\leftarrow 1$;
while SATn is infeasible do
       $n\leftarrow n+1$;
$\{B_{k}\}_{k=1}^{|\mathrm{AP}|}\leftarrow$ SATn solution ;
$\mathcal{G}^{\mathrm{learned}}\leftarrow$ Construct_RM_model($\{B_{k}\}_{k=1}^{|\mathrm{AP}|}$);
$(\mathcal{M}/r)^{\mathrm{learned}}\leftarrow\mathcal{G}^{\mathrm{learned}}\times\mathcal{M}_{L}$;
foreach $\sigma\in(\mathsf{AP})^{l}$ do
       foreach $s\in\mathrm{Reach}(\sigma)$ do
             Define product policy:
$\pi_{\mathrm{Prod}}^{\mathrm{learned}}(a|s,\delta_{\mathbf{u}}^{\mathrm{learned},*}(u_{I},\sigma))\leftarrow\pi_{\mathrm{PTP}}^{\mathrm{true},l}(a|s,\sigma)$
$\delta_{\mathbf{r}}\leftarrow$ IRL_to_extract_reward($\pi_{\mathrm{Prod}}^{\mathrm{learned}},(\mathcal{M}/r)^{\mathrm{learned}}$)
return $\mathcal{R}^{\mathrm{learned}}=(\mathcal{G}^{\mathrm{learned}},\delta_{\mathbf{r}})$;
@@
III-C Proof of Correctness
Let $\mathcal{G}^{\mathrm{learned}}$ be the reward machine model extracted from the SAT solution, with $\delta_{\mathbf{u}}^{\mathrm{learned}}$ being the associated transition function. The first property of our SAT solution is that it is consistent with any fixed depth of the prefix tree policy. We formalize this in the result below.
Proposition 1
Given $\mathcal{M}_{L}$, the depth-$l$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l}$ of the true prefix policy $\pi_{\mathrm{PTP}}^{\mathrm{true}}$, and an upper bound $u_{\mathrm{max}}$ on the number of nodes of the underlying reward machine, let $\mathcal{G}^{\mathrm{learned}}$ be the output of our SAT algorithm, and define $\pi_{\mathrm{PTP}}^{\mathrm{learned}}$ to be the (infinite depth) prefix tree policy induced by $\mathcal{G}^{\mathrm{learned}}$. Then, the learned and the true prefix tree policies have the same depth-$l$ restriction, i.e., $\pi_{\mathrm{PTP}}^{\mathrm{learned},l}=\pi_{\mathrm{PTP}}^{\mathrm{true},l}$.
Proof:
See Appendix A-A. ∎
While Proposition 1 represents a desirable property of our algorithm, being consistent with the first $l$ layers of the true prefix tree policy is not sufficient in general, given that our problem in Section II-D requires consistency with the infinite depth true prefix tree policy. This is potentially problematic if the agent demonstrates unseen changes in its policy beyond the first $l$ layers of the prefix tree policy. At the same time, it is not possible to run our algorithm with an infinite depth prefix tree policy, as the algorithm will never terminate. Fortunately, we can show that if the $l$ is large enough, then additional negative examples will not change the satisfying assignments of the SAT problem beyond $l$ layers.
Proposition 2
@@ III-C Proof of Correctness
Note that, in practice, a depth-$l$ restriction where $l\ll|\mathcal{S}|u_{\mathrm{max}}^{2}$ can be sufficient to find a reward machine that is policy equivalent to the true one if all the solutions of the corresponding SAT problem are policy-equivalent to each other (e.g., they correspond to the same reward machine up to renaming of nodes). This will be further illustrated in the examples section.
IV Examples
For all our examples, we use $\gamma=0.9$ and $\lambda=1.0$ when solving Problem (3) to generate the data for our learning problem. The implementation is done in Python, and the Z3 library [30] is used for solving the SAT problems. Table I summarizes the key parameters related to the problem setup and solution performance for the examples considered in this section. The code is available at github.com/mlshehab/learning_reward_machines
@@ IV Examples
Our first example is on the patrol task of Figure 1. We are given a depth-$6$ restriction of a prefix tree policy $\pi_{\mathrm{PTP}}$, induced by the solution of Problem (3). To enumerate all the satisfying assignments of the SAT problem, we add a constraint that the next solution should be different every time our SAT algorithm finds a solution. We end up with a total of $6$ solutions, which are all the possible renamings of the true reward machine, meaning that the true reward machine is learned up-to-renaming. We also show how the non-stuttering condition of Section III-A4 helps reduce the size of the negative example set, yet still recovering the true reward machine model. While some reduction in the SAT solver time is achieved, the drastic gain is in the time required to encode all the negative examples into the SAT solver, making the overall procedure orders of magnitude faster. For the remaining examples, non-stuttering is assumed.
Depth $|\mathcal{S}|$ $|\mathcal{A}|$ $|\mathcal{E}^{-}|$ SAT time (sec)
Ex.1 w/ (9) 6 16 4 7264 0.81
Ex.1 w/o (9) 6 16 4 168341 6.97
Ex.2 9 9 4 241435 2.859
Ex.3 9 60 9 31656 0.558
Ex.4 8 60 9 24763 0.483
TABLE I: Solution statistics for the examples.
@@
For our second example, we add a hallway between the rooms, as shown in Figure 2. This is a $3\times 3$ grid world, where the corresponding label of each room is shown. The reward machine is kept the same, and the atomic proposition $\mathrm{H}$ is added as a self-loop to all the nodes. With this added hallway, longer atomic proposition prefixes are required to reach all nodes of the reward machine, showing how the underlying MDP affects the required depth for learning a reward machine. For example, the shortest atomic proposition that can reach $u_{3}$ now is $\sigma=\mathrm{AHBHC}$ instead of $\mathrm{ABC}$ as in the previous example. With a depth $9$ prefix tree policy, the reward machine is again learned up-to-renaming.
Figure 2: Hallway Grid World.

### v2

@@ 2.4 Problem Statement
    Does there always exist a depth-$l^{*}$ such that, given the labeled MDP model $\mathcal{M}_{L}$, a bound $u_{\mathrm{max}}$ on the number of nodes of the underlying reward machine, and the depth-$l^{*}$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l^{*}}$ of the true prefix tree policy, it is possible to learn a reward machine that is policy-equivalent to the underlying one?
@@
In what follows, we first provide an algorithm that takes the labeled MDP model $\mathcal{M}_{L}$, the depth-$l$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l}$ of the prefix tree policy $\pi_{\mathrm{PTP}}^{\mathrm{true}}$ for some arbitrary $l$, and the bound $u_{\mathrm{max}}$, and computes a reward machine that induces a prefix tree policy $\pi_{\mathrm{PTP}}^{\mathrm{learned}}$ with the same depth-$l$ restriction, i.e., $\pi_{\mathrm{PTP}}^{\mathrm{learned},l}=\pi_{\mathrm{PTP}}^{\mathrm{true},l}$. Then, we prove the existence of a sufficient depth-$l^{*}$ in (P1), for which this algorithm solves problem (P2). We provide an upper bound on $l^{*}$ in terms of the number of states in the MDP and the number of nodes in the RM.
3 Methodology
3.1 SAT Encoding
@@ 3.1.1 Determinism
3.1.2 Full Specification
This constraint, also known as being input-enabled (Hungar et al., 2003), ensures that all labels generate valid transitions at all nodes. The corresponding Boolean constraints are:
$\forall i,\forall k,\exists j\text{ such that }b_{ikj}=1.$ (6)
We can combine the conditions of Sections 3.1.1 and 3.1.2 into one condition on each $B_{k}$ enforcing that each row has exactly one entry with value $1$.
3.1.3 Negative Examples
@@
Proof.
It follows from (3). ∎
@@
$\mathcal{E}^{-}_{l}=\{\{\sigma,\sigma^{\prime}\}\mid\pi_{\mathrm{PTP}}^{l}(a|s,\sigma)\neq\pi_{\mathrm{PTP}}^{l}(a|s,\sigma^{\prime})\text{ forsome }s,a\}.$ (7)
Let $\sigma=k_{1}k_{2}\cdots k_{l}$ and $\sigma^{\prime}=k_{1}^{\prime}k_{2}^{\prime}\cdots k_{m}^{\prime}$ be two propositional prefixes that lead to different policies in the same state, therefore $\{\sigma,\sigma^{\prime}\}\in\mathcal{E}^{-}_{l}$. We encode the condition given by Lemma 1 into Boolean constraints as:
$(B_{k_{l}}^{\intercal}B_{k_{l-1}}^{\intercal}\cdots B_{k_{1}}^{\intercal}e_{1})\bigwedge(B_{k_{m}^{\prime}}^{\intercal}B_{k_{m-1}^{\prime}}^{\intercal}\cdots B_{k_{1}^{\prime}}^{\intercal}e_{1})=$ $\begin{bmatrix}0&0&\cdots&0\end{bmatrix}^{\intercal},$ (8)
where $\bigwedge$ is the element-wise AND operator and $e_{1}\triangleq[1,0,\cdots,0]^{\intercal}$ indicates that the paths start from the initial node. Our algorithm adds the Boolean constraint given in (8) for each element of $\mathcal{E}^{-}$. The significance of encoding negative examples is that it eliminates the learning of trivial reward machines. In particular, our method never learns a trivial one-state reward machine with all self-transitions (Icarte et al., 2023) as long as there is at least one negative example in our prefix tree policy.
3.1.4 Non-Stuttering
A reward machine is said to be non-stuttering if when a proposition transitions into a node, that same proposition can not transition out of the node, i.e., for all $(u,a,u^{\prime})\in\mathcal{U}\times\mathrm{AP}\times\mathcal{U}$: $\delta_{\mathbf{u}}(u,a)=u^{\prime}\implies\delta_{\mathbf{u}}(u^{\prime},a)=u^{\prime}$. This is related to multi-stage tasks where the particular duration spent on a subtask (i.e., satisfying a given atomic proposition) is not important (Baier and Katoen, 2008). When the reward machine is a priori known to be non-stuttering, this extra condition can be included in the SAT problem. The main significance of this condition is trace-compression (Icarte et al., 2023), by which we can reduce the number of negative examples in $\mathcal{E}^{-}$ by only keeping the shortest negative examples among the equivalent ones. In this case, two negative examples are equivalent if between two pairs, the corresponding label sequences differ only by the same proposition repeated consecutively more than one time. We encode non-stuttering into Boolean constraints as follows:
@@ 3.1.4 Non-Stuttering
The utility of this constraint is demonstrated empirically in Section 5.1.
3.2 Algorithm
@@
The last step is finding the numerical values of rewards that render the product policy $\pi_{\mathrm{Prod}}^{\mathrm{learned}}$ optimal for the product MDP model $(\mathcal{M}/r)^{\mathrm{learned}}$. This is a standard IRL problem without reward machines where the special structure of the rewards on the product can be represented as features. We solve this step using the method developed in (Shehab et al., 2024). Featurization (see, (Shehab et al., 2024, Section 4)) is used to enforce that the reward function of the product can be written as $r((s,u),a,(s^{\prime},u^{\prime}))=\delta_{\mathbf{r}}(u,L(s^{\prime}))$, with $u^{\prime}=\delta_{\mathbf{u}}(u,L(s^{\prime}))$. This IRL method gives us the corresponding output function $\delta_{\mathbf{r}}$ of the reward machine.
Remark 1.
Although we constrain the output function of the reward machine to be of the form $\delta_{r}:\mathcal{U}\times\mathrm{AP}\to\mathbb{R}$ (leading to what is commonly known as simple reward machines (Icarte et al., 2018)), this does not limit the generality of our framework. The same procedure remains applicable in the case of a dense output function of the form $\delta_{r}:\mathcal{U}\times\mathrm{AP}\to[\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}]$. We focus on the simple output function formulation in this paper to simplify both the presentation and the interpretation of the resulting reward machines.
@@
1
[…]
3while SATn is infeasible do
4       $n\leftarrow n+1$
5
6Construct $\mathcal{E}_{l}^{-}$ using (7)
7$\{B_{k}\}_{k=1}^{|\mathrm{AP}|}\leftarrow$ SATn solution
[…]
9$(\mathcal{M}/r)^{\mathrm{learned}}\leftarrow\mathcal{G}^{\mathrm{learned}}\times\mathcal{M}_{L}$
10foreach $\sigma\in(\mathsf{AP})^{l}$ do
11       foreach $s\in\mathrm{Reach}(\sigma)$ do
12             Define product policy: $\pi_{\mathrm{Prod}}^{\mathrm{learned}}(a|s,\delta_{\mathbf{u}}^{\mathrm{learned},*}(u_{I},\sigma))\leftarrow\pi_{\mathrm{PTP}}^{\mathrm{true},l}(a|s,\sigma)$
13
14
[…]
return $\mathcal{R}^{\mathrm{learned}}=(\mathcal{G}^{\mathrm{learned}},\delta_{\mathbf{r}})$
@@ 4       $n\leftarrow n+1$
Remark 2.
It is important to emphasize that up to line 7 in Algorithm 1, neither the optimality of the prefix tree policy nor the transition kernel $\mathcal{P}$ of the MDP is used. That is, our method learns a reward machine model $\mathcal{G}^{\mathrm{learned}}$ in a model-free fashion as long as the prefix-tree policy is induced by a reward machine. The optimality with respect to the MaxEnt objective in (1) and the transition kernel only comes into play to extract the numerical reward values in lines 8-11.
3.3 Proof of Correctness
Let $\mathcal{G}^{\mathrm{learned}}$ be the reward machine model extracted from the SAT solution, with $\delta_{\mathbf{u}}^{\mathrm{learned}}$ being the associated transition function. The first property of our SAT solution is that it is consistent with any fixed depth-of the prefix tree policy. We formalize this in the result below.
Proposition 1.
Given the labeled MDP model $\mathcal{M}_{L}$, the depth-$l$ restriction $\pi_{\mathrm{PTP}}^{\mathrm{true},l}$ of the true prefix policy $\pi_{\mathrm{PTP}}^{\mathrm{true}}$, and an upper bound $u_{\mathrm{max}}$ on the number of nodes of the underlying reward machine, let $\mathcal{G}^{\mathrm{learned}}$ be the output of our SAT problem, and define $\pi_{\mathrm{PTP}}^{\mathrm{learned}}$ to be the (infinite depth) prefix tree policy induced by $\mathcal{G}^{\mathrm{learned}}$. Then, the learned and the true prefix tree policies have the same depth-$l$ restriction, i.e., $\pi_{\mathrm{PTP}}^{\mathrm{learned},l}=\pi_{\mathrm{PTP}}^{\mathrm{true},l}$.
Proof.
See Appendix A.1. ∎
While Proposition 1 represents a desirable property of our algorithm, being consistent with the depth-$l$ restriction of the true prefix tree policy is in general not sufficient to be consistent with the true (infinite-depth) prefix tree policy (as required in Problem (P1) from Section 2.4). This is potentially problematic if the agent demonstrates unseen changes in its policy for prefixes longer than $l$. At the same time, it is not possible to run our algorithm with the unrestricted prefix tree policy $\pi_{\mathrm{PTP}}^{\mathrm{true}}$ because it would lead to an infinite number of negative examples, i.e., $|\mathcal{E}^{-}_{\infty}|=\infty$. Consequently, the algorithm would not terminate. Fortunately, we can show that if $l$ is large enough, then increasing $l$ will not change the satisfying assignments of the SAT problem.
Proposition 2.
@@ 3.3 Proof of Correctness
Note that, in practice, a depth-$l$ restriction where $l\ll|\mathcal{S}|u_{\mathrm{max}}^{2}$ can be sufficient to find a reward machine that is policy equivalent to the true one if all the solutions of the corresponding SAT problem are policy-equivalent to each other (e.g., they correspond to the same reward machine up to renaming of nodes). This will be further illustrated in the experiments of Section 5.1.
4 Learning From Demonstrations
In this section, we present how our method can be applied when the optimal policy is known only through a set of optimal demonstrations $\mathfrak{D}=\{(\sigma_{i},s_{i},a_{i})\}_{i=1}^{|\mathfrak{D}|}$. Hence, the depth-$l$ restriction of the true prefix tree policy is unknown. Consequently, we construct the following consistent unbiased estimate:
$\hat{\pi}_{\mathrm{PTP}}(a|s,\sigma)=\frac{\sum_{i=1}^{|\mathfrak{D}|}\mathbb{I}(\sigma_{i}=\sigma,s_{i}=s,a_{i}=a)}{\sum_{i=1}^{|\mathfrak{D}|}\mathbb{I}(\sigma_{i}=\sigma,s_{i}=s)}.$ (11)
However, the fact that we only have access to an approximation of $\pi_{\mathrm{PTP}}$ leads to two challenges. First, using $\hat{\pi}_{\mathrm{PTP}}$ to directly construct the set of negative examples $\mathcal{E}^{-}_{l}$ will lead to many pairs $\{\sigma,\sigma^{\prime}\}$ being incorrectly considered as negative examples (see equation (7)). Second, the SAT problem is not robust to incorrect negative examples. In the next subsections, we address these two problems.
4.1 Estimating Negative Examples
To limit the inclusion of incorrect negative examples in $\mathcal{E}^{-}_{l}$, a pair of prefixes $\{\sigma_{1},\sigma_{2}\}$ will be considered as a negative example only if we have high confidence that they correspond to negative examples given the true prefix tree policy. In particular, let $\sigma_{1},\sigma_{2}$ be two prefixes from $\mathfrak{D}$ that we want to compare and let $s\in\mathcal{S}$ be a state. For $j\in\{1,2\}$, let $n_{j}$ be the number of visitations to the pair $(s,\sigma_{j})$, i.e., the denominator in (11). Note that $\pi_{\mathrm{PTP}}(a|s,\sigma_{j})$ is a categorical distribution of which we have a sample estimate for. Hence, it follows from (Weissman et al., 2003) that for all $\epsilon>0$:
$\mathbb{P}(\|\hat{\pi}_{\mathrm{PTP}}(a|s,\sigma_{j})-\pi_{\mathrm{PTP}}(a|s,\sigma_{j})\|_{1}\leq\epsilon)$ $\geq 1-\delta_{j},$ (12)
with $\delta_{j}\triangleq(2^{|\mathcal{A}|}-2)e^{-\frac{1}{2}n_{j}\epsilon^{2}}$. We pick $\epsilon=\frac{1}{2}\|\hat{\pi}_{\mathrm{PTP}}(a|s,\sigma_{1})-\hat{\pi}_{\mathrm{PTP}}(a|s,\sigma_{2})\|_{1}$. Using a union bound, the probability of the two confidence intervals described in (12) (for $j\in\{1,2\}$) to overlap is lower bounded by $1-\delta_{1}-\delta_{2}$. Consequently, we consider the pair $\{\sigma_{1},\sigma_{2}\}$ as being a negative example if there is a state $s$ such that $1-\delta_{1}-\delta_{2}\geq 1-\alpha$, where $\alpha$ is a user-defined parameter.
4.2 MAX-SAT: SAT with Robustness to Incorrect Negative Examples
Even when the set of negative examples is constructed from pairs of prefixes for which we have a high confidence of being true negative examples, some pairs may still be mislabeled as negative examples. We refer to those as false positives. To deal with these false positive, we implement a weighted MAX-SAT (Manquinho et al., 2009; Biere et al., 2009) variant of the SAT problem. Concretely, the weighted MAX-SAT problem consists of finding a Boolean assignment of the variables $b_{ikj}$ such that (i) constraints (5), (6) and (9) hold, i.e., the recovered RM is deterministic, non-stuttering, and fully specified; and (ii) the constraint (8) holds for a maximum number of pairs $\{\sigma,\sigma^{\prime}\}$ in $\mathcal{E}^{-}_{l}$. In other words, (5), (6) and (9) are hard constraints whereas the constraints coming from (8) are soft constraints. This approach allows for robustness to misclassifications of negative examples. We empirically demonstrate in Section 5.3 that this weighted MAX-SAT formulation consistently infers the smallest set of consistent reward machine models, even in the presence of false positives.
5 Experiments
To demonstrate the generality and efficiency of our approach, we apply it to a diverse set of domains, from classical grid-based MDPs to a continuous robotic control task and a real-world biological navigation dataset. In every experiment, we fix the discount factor to $\gamma=0.99$ and the regularization weight to $\lambda=1.0$ when solving Problem (2), both for generating demonstration traces and for reward recovery. Our code is implemented in Python, and the Z3 library (De Moura and Bjørner, 2008) is used for solving the SAT and weighted MAX-SAT problems. To enumerate all the satisfying assignments of the SAT problem, we add a constraint that the next solution should be different every time our SAT solver finds a solution. We will release the complete implementation on GitHub, upon acceptance to adhere with the double-blind review process.
5.1 Tabular GridWorld MDPs
@@
Our first experiment (patrol) is on the running-example patrol task of Figure 1. By setting $u_{\mathrm{max}}=4$, our bound from Theorem 1 is $256$. However, with only a depth-$6$ prefix tree policy, we end up with a total of $\mathbf{6}$ solutions. These are all the possible renamings of the true reward machine (see Figure 1(b)), meaning that the true reward machine is learned up-to-renaming with a depth-$6$ prefix tree policy. Table 1 summarizes the findings, where we also show how the non-stuttering condition of Section 3.1.4 helps reduce the size of the negative example set, yet still recovering the true reward machine model. While some reduction in the SAT solver time is achieved, the drastic gain is in the time required to encode all the negative examples into the SAT solver, making the overall procedure orders of magnitude faster. For the remaining experiments, non-stuttering is assumed.
Task Depth $|\mathcal{S}|$ $|\mathcal{A}|$ $|\mathcal{E}^{-}|$ SAT time (sec)
patrol with (9) 6 16 4 7264 0.81
patrol without (9) 6 16 4 168341 6.97
patrol-hallway 9 9 4 241435 2.859
Table 1: Solution statistics for the tabular GridWorld MDP.
@@
For our second experiment (patrol-hallway), we add a hallway between the rooms, as shown in Figure 1(c). This is a $3\times 3$ grid world, where the corresponding label of each room is shown. The reward machine is kept the same, and the atomic proposition $\mathrm{H}$ is added as a self-loop to all the nodes. With this added hallway, longer atomic proposition prefixes are required to reach all nodes of the reward machine, showing how the underlying MDP affects the required depth for learning a reward machine. For example, the shortest atomic proposition that can reach $u_{3}$ is now $\sigma=\mathrm{AHBHC}$ instead of $\mathrm{ABC}$ as in the previous example. With a depth-$9$ prefix tree policy, the reward machine is again learned up-to-renaming.
5.2 Tabular BlockWorld MDPs